OpenStudy (anonymous):

Amanda tells you that when variables are in the denominator, the equation becomes unsolvable. "There is a value for x that makes the denominator zero, and you can't divide by zero," Amanda explains. Using complete sentences, demonstrate to Amanda how the equation is still solvable.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

I made an equation for it such as 3x^2+x/x

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to solve that equation?

OpenStudy (anonymous):

I did and got the answer 3x+1 but how would this respond to this question

jimthompson5910 (jim_thompson5910):

well if you had something like \[\Large \frac{3x^2+x}{x-1} = 0\] then multiplying both sides by x-1 gives you \[\Large \frac{3x^2+x}{x-1} = 0\] \[\Large 3x^2+x = 0(x-1)\] \[\Large 3x^2+x = 0\]

jimthompson5910 (jim_thompson5910):

what do you get when you solve \[\Large 3x^2+x = 0\]

OpenStudy (anonymous):

x=0 or x=1/3

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

x = 0 is correct

OpenStudy (anonymous):

-1/3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that proves just because you have a variable in the denominator, it doesn't mean the equation isn't solveable

OpenStudy (anonymous):

oh ok! that would be the solution?

jimthompson5910 (jim_thompson5910):

those two x values, yes

jimthompson5910 (jim_thompson5910):

their existence shows us that the initial claim is false